# Classification of nilpotent almost abelian Lie groups admitting left-invariant complex or symplectic structures

### v1

@@ 3. Classification of nilpotent almost abelian Lie algebras admitting complex structures
Let $\mathfrak{g}_{A}={\mathbb{R}}e_{0}\ltimes_{A}{\mathbb{R}}^{2n-1}$ be a $k$-step nilpotent almost abelian Lie algebra. If $\mathfrak{g}_{A}$ admits a complex structure then $k\leq n$.
@@
If $\mathfrak{g}_{A}$ admits a complex structure, then $A$ is conjugate to a matrix $C$ of the form (8) with $a=0$ and $J=J_{n-1}$ as in (5) (see Remarks 2.2 and 3.4). In particular, $B\in M_{J_{n-1}}(2(n-1),{\mathbb{R}})={\rm Im}\,\varphi$, with $\varphi:M(n-1,{\mathbb{C}})\to M_{J_{n-1}}(2(n-1),{\mathbb{R}})$ as in (4). Since $\mathfrak{g}_{A}$ is $k$-step nilpotent, $C^{k}=0$ and $C^{k-1}\neq 0$, which implies that $B^{k}=0$.
If $B^{k-1}\neq 0$, then $\varphi^{-1}(B)\in M(n-1,{\mathbb{C}})$ satisfies $(\varphi^{-1}(B))^{k}=0$ and $(\varphi^{-1}(B))^{k-1}\neq 0$, that is, the minimal polynomial of $\varphi^{-1}(B)$ is $x^{k}$, therefore, $k\leq n-1$.
If $B^{k-1}=0$, then $B^{k-2}\neq 0$ since $C^{k-1}\neq 0$. In this case we have that $(\varphi^{-1}(B))^{k-1}=0$ and $(\varphi^{-1}(B))^{k-2}\neq 0$, and the minimal polynomial of $\varphi^{-1}(B)$ is $x^{k-1}$, hence $k-1\leq n-1$, and the corollary follows. ∎
@@
with $B=0_{2k+2}\oplus\mathcal{J}_{2}^{\oplus(p-1)}\in M_{J}(2(n-1),{\mathbb{R}})$, for $J=\tilde{J}\oplus J_{2l}$, where $\tilde{J}$ is an arbitrary $2(k+1)\times 2(k+1)$ matrix such that $(\tilde{J})^{2}=-I_{2(k+1)}$. ∎
@@ 2. ${\rm(ii)}$
3.1. Classification Theorem (Complex Case)
Theorem 3.10 below gives the classification of the nilpotent almost abelian Lie algebras admitting a complex structure (compare with [4,  Appendix A]).
The next proposition will be useful to prove Theorem 3.10.
Proposition 3.9.
Let $B\in M(2m,{\mathbb{R}})$ given by
$B=\mathcal{J}_{m_{1}}^{\oplus q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}^{\oplus q_{r}}\oplus 0_{s},$
with $m_{1}>\cdots>m_{r}\geq 2$, $q_{i}\geq 1$ for all $i$, $s\geq 0$ and $\sum_{i=1}^{r}q_{i}m_{i}+s=2m$. Then, $B$ is conjugate in $GL(2m,{\mathbb{R}})$ to a matrix in $M_{J_{m}}(2m,{\mathbb{R}})$ if and only if $s$ is even and $q_{i}$ is even for all $i$.
Proof.
Suppose that $B$ is conjugate to $B^{\prime}\in M_{J_{m}}(2m,{\mathbb{R}})$. Via the isomorphism $\varphi$ given in (4), if the Jordan normal form of $\varphi^{-1}(B^{\prime})$ is given by:
$\mathcal{J}_{d_{1}}\oplus\cdots\oplus\mathcal{J}_{d_{k}}\oplus 0_{t},\qquad d_{1}\geq\cdots\geq d_{k}\geq 2,\quad t\geq 0,$
then, by uniqueness, $B$ has the following Jordan normal form:
$\mathcal{J}^{\oplus 2}_{d_{1}}\oplus\cdots\oplus\mathcal{J}^{\oplus 2}_{d_{k}}\oplus 0_{2t},\qquad d_{1}\geq\cdots\geq d_{k}\geq 2,\quad t\geq 0.$
In particular, each Jordan block in the Jordan normal form of $B$ appears an even number of times, that is, $s$ is even and $q_{i}$ is even for all $i$.
Conversely, assume that $s=2t$ and $q_{i}=2p_{i}$, for all $i$. We consider $B_{1}=\mathcal{J}_{m_{1}}^{\oplus p_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}^{\oplus p_{r}}\oplus 0_{t}$. Then $B$ is conjugate to $B_{1}\oplus B_{1}\in M_{J_{m}}(2m,{\mathbb{R}})$. This concludes the proof. ∎
@@ 3.1. Classification Theorem (Complex Case)
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{2n-1}$ be the $2n$-dimensional nilpotent almost abelian Lie algebra with
$M=\mathcal{J}_{n_{1}}^{\oplus p_{1}}\oplus\cdots\oplus\mathcal{J}_{n_{k}}^{\oplus p_{k}}\oplus 0_{t},\quad n_{1}>\cdots>n_{k}\geq 2,\;\,p_{i}\geq 1\text{for all }i,\;t\geq 0,$
where $\sum_{i=1}^{k}p_{i}n_{i}+t=2n-1$. Then $\mathfrak{g}_{M}$ admits a complex structure if and only if one of the following conditions holds:
@@ 3. ${\rm(iii)}$
Assume that $\mathfrak{g}_{M}$ admits a complex structure, then by Remarks 3.4 and 3.2, $M$ is conjugate to a matrix
@@
Then, by Proposition 3.9. $B$ is conjugate to $B^{\prime}=\mathcal{J}_{m_{1}}^{\oplus 2q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}^{\oplus 2q_{r}}\oplus 0_{2s}$, where $m_{1}>\dots>m_{r}\geq 2$. Hence, $M$ is conjugate to
@@
For the converse, by Remark 3.4 it is enough to show (case by case) that $M$ is conjugated to a matrix $C$ as in  (8).
@@
$C=\begin{pmatrix}0&0\\ \epsilon_{1}&B\end{pmatrix},\quad\text{ where }\quad B=0_{2(s+1)}\oplus\mathcal{J}_{n_{1}}^{\oplus 2q_{1}}\oplus\cdots\oplus\mathcal{J}_{2}^{\oplus 2q_{k}}.$
@@
4. Classification of nilpotent almost abelian Lie algebras admitting symplectic structures

### v2

@@ 4.2. The nilpotent case
Let $\mathfrak{g}_{A}={\mathbb{R}}e_{0}\ltimes_{A}{\mathbb{R}}^{2n-1}$ be a $k$-step nilpotent almost abelian Lie algebra. If $\mathfrak{g}_{A}$ admits a complex structure, then $k\leq n$.
@@
If $\mathfrak{g}_{A}$ admits a complex structure, then $A$ is conjugate to a matrix $C$ of the form (12) with $a=0$ and $J=J_{n-1}$ as in (10) (see Remark 2.2 and Remark 4.4). In particular, $B\in M_{J_{n-1}}(2(n-1),{\mathbb{R}})={\rm Im}\,\varphi$, with $\varphi:M(n-1,{\mathbb{C}})\to M_{J_{n-1}}(2(n-1),{\mathbb{R}})$ as in (9). Since $\mathfrak{g}_{A}$ is $k$-step nilpotent, $C^{k}=0$ and $C^{k-1}\neq 0$, which implies that $B^{k}=0$.
If $B^{k-1}\neq 0$, then $\varphi^{-1}(B)\in M(n-1,{\mathbb{C}})$ satisfies $(\varphi^{-1}(B))^{k}=0$ and $(\varphi^{-1}(B))^{k-1}\neq 0$, that is, the minimal polynomial of $\varphi^{-1}(B)$ is $x^{k}$. Therefore, $k\leq n-1$.
If $B^{k-1}=0$, then $B^{k-2}\neq 0$ since $C^{k-1}\neq 0$. In this case we have that $(\varphi^{-1}(B))^{k-1}=0$ and $(\varphi^{-1}(B))^{k-2}\neq 0$. Then, the minimal polynomial of $\varphi^{-1}(B)$ is $x^{k-1}$, hence $k-1\leq n-1$, and the corollary follows. ∎
@@
with $B=0_{2k+2}\oplus\mathcal{J}_{2}^{\oplus(p-1)}\in M_{J}(2(n-1),{\mathbb{R}})$, for $J=\tilde{J}\oplus J_{2l}$, where $\tilde{J}$ is an arbitrary $2(k+1)\times 2(k+1)$ matrix such that $\tilde{J}^{2}=-I_{2(k+1)}$. ∎
@@ 2. ${\rm(ii)}$
Given a nilpotent matrix $M$, in the following theorem we give necessary and sufficient conditions on the tuple $\Lambda(M)$ to obtain pairwise non-isomorphic nilpotent almost abelian Lie algebras $\mathfrak{g}_{M}$ admitting complex structures (compare with [2,  Appendix A]). We exclude the zero matrix $M$, since when $M=0$, $\mathfrak{g}_{M}$ is the abelian Lie algebra, which always admits a complex structure.
@@
Let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{2n-1}$ be the $2n$-dimensional nilpotent almost abelian Lie algebra with $\Lambda(M)=(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t)$. Then $\mathfrak{g}_{M}$ admits a complex structure if and only if one of the following conditions holds:
@@ 3. ${\rm(iii)}$
Assume that $\mathfrak{g}_{M}$ admits a complex structure, then by Remark 4.4 and Remark 4.2, $M$ is conjugate to a matrix
@@
Then, either $B=0$ and $v\neq 0$, or $B\neq 0$. If $B=0$, Corollary 3.7 implies that $C$, hence $M$, is conjugate to $\mathcal{J}_{2}\oplus 0_{2n-3}$. Therefore, $M$ satisfies condition (ii).
If $B\neq 0$, by Proposition 4.5 $B$ is conjugate to $B^{\prime}=\mathcal{J}_{m_{1}}^{\oplus 2q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}^{\oplus 2q_{r}}\oplus 0_{2s}$, where $m_{1}>\dots>m_{r}\geq 2$. Hence, $M$ is conjugate to
@@
For the converse, by Remark 4.4 it is enough to show (case by case) that $M$ is conjugate to a matrix $C$ as in  (12).
@@
$C=\begin{pmatrix}0&0\\ \epsilon_{1}&B\end{pmatrix},\quad\text{ where }\quad B=0_{2(s+1)}\oplus\mathcal{J}_{n_{1}}^{\oplus 2q_{1}}\oplus\cdots\oplus\mathcal{J}_{2}^{\oplus 2q_{k}}\;\text{or}\;B=0_{2(n-1)}.$
@@
For each $n\in{\mathbb{N}}$, let
[…]
Consider the bijection $\tilde{\Lambda}$ defined in (8) for $b=0$ and $m=2n-1$. As a consequence of Lemma 3.3 and Theorem 4.10, we obtain the following classification result.
Corollary 4.11.
Let $n\in{\mathbb{N}}$ and let $\mathfrak{g}_{M}={\mathbb{R}}e_{0}\ltimes_{M}{\mathbb{R}}^{2n-1}$ be a $2n$-dimensional nilpotent almost abelian Lie algebra. Then $\mathfrak{g}_{M}$ admits a complex structure if and only if $\Lambda(M)\in{\bigwedge}^{c}_{2n-1}$. Moreover, the isomorphism classes of $2n$-dimensional nilpotent almost abelian Lie algebras admitting a complex structure are parametrized by the set $\bigwedge^{c}_{2n-1}$.
Example 4.12.
We obtain that the set ${\bigwedge}^{c}_{7}$ consists of the following tuples:
$(4,3;1,1;0),\,(3;2;1),\,(3,2;1,1;2),\,(2;3;1),\,(2;2;3),\,(2;1;5),\,(7).$
Therefore, there are 7 isomorphism classes of 8-dimensional nilpotent almost abelian Lie algebras admitting complex structures. The set ${\bigwedge}_{7\,}-{\bigwedge}^{c}_{7}$ contains 8 tuples, that is, there are 8 isomorphism classes of 8-dimensional nilpotent almost abelian Lie algebras not admitting complex structures. For instance, since $(7;1;0)\notin{\bigwedge}^{c}_{7}$, the corresponding nilpotent almost abelian Lie algebra does not admit a complex structure.
Example 4.13.
The set ${\bigwedge}^{c}_{9}$ is given by
$\begin{array}[]{llllll}(5,4;1,1,0),&(4;2;1),&(4,3;1,1;2),&(3,2:2,1;1),&(3;2;3),&(3,2;1,3;0),\\ (3,2;1,1;4),&(2;4;1),&(2;3;3),&(2;2;5),&(2;,1,7),&(9).\end{array}$
This gives that there are 12 isomorphism classes of 10-dimensional nilpotent almost abelian Lie algebras admitting complex structures. It can be shown that the set ${\bigwedge}_{9\,}-{\bigwedge}^{c}_{9}$ contains 18 tuples, that is, there are 18 isomorphism classes of 10-dimensional nilpotent almost abelian Lie algebras not admitting complex structures. For instance, $(9;1;0)\notin{\bigwedge}^{c}_{9}$.
Example 4.14.
When $M=\mathcal{J}_{2n-1}$, with $n\geq 2$, the Lie algebra $\mathfrak{g}_{M}$ is a filiform Lie algebra. Since $(2n-1;1;0)\notin{\bigwedge}^{c}_{2n-1}$, $\mathfrak{g}_{M}$ does not admit a complex structure, a fact that was already known (see [28]).
4.3. The general case
We start by proving the following result, which will be needed in Theorem 4.16 and Theorem 4.17.
Proposition 4.15.
Let $M\in M(2n-1,{\mathbb{R}})$ and assume that $\mathfrak{g}_{M}$ admits a complex structure, then there exists a unique $a\in\text{spec}_{\mathbb{R}}(M)$ such that $m_{a}$ is odd. In particular, $M$ is conjugate to either $M(a)$ or $M(a)\oplus Q$, where $Q\in M(2n-1-m_{a},{\mathbb{R}})$ and $a\not\in\emph{spec}\,(Q)$, depending on whether $m_{a}=2n-1$ or $m_{a}<2n-1$.
Proof.
By Remark 4.4 $M$ is conjugate to $C=\begin{pmatrix}a&0\\ v&B\end{pmatrix}$, where $a\in{\mathbb{R}}$, $v\in{\mathbb{R}}^{2n-2}$ and $B\in M_{J}(2n-2,{\mathbb{R}})$, for some $J$. Since $M$ and $C$ are conjugate and $a$ is an eigenvalue of $C$, then $a$ is an eigenvalue of $M$. By Proposition 4.5, $a$ is the unique real eigenvalue of $M$ such that $m_{a}$ is odd. The last sentence of the statement follows directly from Lemma 3.1. ∎
In view of Proposition 4.15 we state the following theorems. In the first one, which is a consequence of Corollary 4.11, we consider the case when $M$ has a unique eigenvalue $a$, with $a\in{\mathbb{R}}$.
Theorem 4.16.
Let $M\in M_{a}(2n-1,{\mathbb{R}})$, with $a\in{\mathbb{R}}$. Then, $\mathfrak{g}_{M}$ admits a complex structure if and only if $\Lambda(M)\in{\bigwedge}^{c}_{2n-1}$.
Proof.
Suppose that $\mathfrak{g}_{M}$ admits a complex structure, then by Remark 4.4, $M$ is conjugate to $C=\begin{pmatrix}a&0\\ v&B\end{pmatrix}$, where $v\in{\mathbb{R}}^{2n-2}$ and $B\in M_{J}(2n-2,{\mathbb{R}})$ for some $J$. In particular, $B\in M_{a}(2n-2,{\mathbb{R}})$. We consider the nilpotent matrix $N=C-aI_{2n-1}=\begin{pmatrix}0&0\\ v&N^{\prime}\end{pmatrix}$, where $N^{\prime}=B-aI_{2n-2}$. Since $N^{\prime}\in M_{J}(2n-2,{\mathbb{R}})$, by Remark 4.4 we obtain that $\mathfrak{g}_{N}$ admits a complex structure. Then, Corollary 4.11 implies that $\Lambda(N)\in{\bigwedge}^{c}_{2n-1}$. Finally, using Remark 3.2 and the fact that $\Lambda(C)=\Lambda(M)$, we obtain that $\Lambda(M)\in{\bigwedge}^{c}_{2n-1}$.
Conversely, suppose that $\Lambda(M)\in{\bigwedge}^{c}_{2n-1}$ and consider as above, the nilpotent matrix $N=M-aI_{2n-1}$, then by Remark 3.2 we have $\Lambda(N)\in{\bigwedge}^{c}_{2n-1}$. Therefore, Corollary 4.11 implies that $\mathfrak{g}_{N}$ admits a complex structure. In particular, $N$ is conjugate to a matrix $\left(\begin{array}[]{cc}0&0\\ v&N^{\prime}\end{array}\right)$, where $N^{\prime}\in M_{J}(2n-2,{\mathbb{R}})$ for some $J$. So, $M$ is conjugate to $\left(\begin{array}[]{cc}a&0\\ v&B\end{array}\right)$, where $B=N^{\prime}+aI_{2n-2}\in M_{J}(2n-2,{\mathbb{R}})$. In consequence, by Remark 4.4, $\mathfrak{g}_{M}$ admits a complex structure. ∎
In the next theorem we consider the case when $M$ has two or more distinct eigenvalues and at least one of them is real.
Theorem 4.17.
Let $M\in M(2n-1,{\mathbb{R}})$ such that $M=M(a)\oplus Q$ where $a\not\in\emph{spec}\,(Q)$, $M(a)\in M(m_{a},{\mathbb{R}})$, with $a\in{\mathbb{R}}$ and $m_{a}$ odd. Then, $\mathfrak{g}_{M}$ admits a complex structure if and only if $\Lambda(M(a))\in{\bigwedge}^{c}_{m_{a}}$ and $Q\in M_{J^{\prime}}(m,{\mathbb{R}})$, for some $J^{\prime}$ such that $(J^{\prime})^{2}=-I_{m}$ with $m=2n-1-m_{a}$.
Proof.
First, assume that $M(a)=aI_{m_{a}}$, then $M=aI_{m_{a}}\oplus Q$ with $m_{a}$ odd. If $m_{a}=1$, then $\mathfrak{g}_{M}$ admits a complex structure if and only if $Q\in M_{J^{\prime}}(m,{\mathbb{R}})$, for some $J^{\prime}$. On the other hand, if $m_{a}>1$ we have that $m_{a}-1>0$ is even and $aI_{m_{a}-1}\oplus Q\in M_{J^{\prime}}(2n-2,{\mathbb{R}})$, for some $J^{\prime}$, and this holds if and only if $Q\in M_{J^{\prime}}(m,{\mathbb{R}})$, for some $J^{\prime}$. Therefore, by Remark 4.4, $\mathfrak{g}_{M}$ admits a complex structure if and only if $Q\in M_{J^{\prime}}(m,{\mathbb{R}})$, for some $J^{\prime}$.
Now, assume that $M(a)\neq aI_{m_{a}}$. If $\mathfrak{g}_{M}$ admits a complex structure, then by Remark 4.4, $M$ is conjugate to $C=\begin{pmatrix}a&0\\ v&B\end{pmatrix}$, where $v\in{\mathbb{R}}^{2n-2}$ and $B\in M_{J}(2n-2,{\mathbb{R}})$ for some $J$. Therefore, by Proposition 4.5, any real eigenvalue of $B$ satisfies condition (13). Since $M(a)\neq aI_{m_{a}}$, we have that $m_{a}>1$. Hence, $a\in\text{spec}_{\mathbb{R}}(B)$ and $B$ is conjugate to $B(a)\oplus Q^{\prime}$, with $a\notin\text{spec}\,(Q^{\prime})$. Then, any real eigenvalue of both, $B(a)$ and $Q^{\prime}$, satisfies condition (13). In particular, by Proposition 4.5, $B(a)\in M_{J_{1}}(m_{a}-1,{\mathbb{R}})$ and $Q^{\prime}\in M_{J_{2}}(m,{\mathbb{R}})$, for some $J_{1}$ and $J_{2}$. Since $a\notin\text{spec}(Q^{\prime})$, Lemma 3.8 implies that $C$ is conjugate to $\left(\begin{array}[]{c|c|c}a&0&0\\ \hline\cr u&B(a)&0\\ \hline\cr 0&0&Q^{\prime}\end{array}\right)$, where $u\in{\mathbb{R}}^{m_{a}-1}$, $B(a)\in M(m_{a}-1,{\mathbb{R}})$, $Q^{\prime}\in M(m,{\mathbb{R}})$.
Let $C_{a}:=\left(\begin{array}[]{cc}a&0\\ u&B(a)\end{array}\right)\in M(m_{a},{\mathbb{R}})$. Since $B(a)\in M_{J_{1}}(m_{a}-1,{\mathbb{R}})$, by Remark 4.4, $\mathfrak{g}_{C_{a}}$ admits a complex structure. Then applying Theorem 4.16 we have $\Lambda(C_{a})\in{\bigwedge}^{c}_{m_{a}}$. Finally, since $M=M(a)\oplus Q$ is conjugate to $C_{a}\oplus Q^{\prime}$, we obtain that $\Lambda(C_{a})=\Lambda(M(a))$ and $Q^{\prime}$ is conjugate to $Q$. Hence, the statement follows.
Conversely, suppose that $\Lambda(M(a))\in{\bigwedge}^{c}_{m_{a}}$, then by Theorem 4.16, we obtain that $\mathfrak{g}_{M(a)}$ admits a complex structure. In particular, $M(a)$ is conjugate to a matrix $\left(\begin{array}[]{cc}a&0\\ v&B\end{array}\right)$, where $B\in M_{J_{1}}(m_{a}-1,{\mathbb{R}})$ for some $J_{1}$. On the other hand, $Q\in M_{J^{\prime}}(m,{\mathbb{R}})$ for some $J^{\prime}$, since any real eigenvalue of $Q$ satisfies condition (13). Finally, by Remark 4.4, $\mathfrak{g}_{M}$ admits a complex structure since $M$ is conjugate to $\left(\begin{array}[]{c|c|c}a&0&0\\ \hline\cr v&B&0\\ \hline\cr 0&0&Q\end{array}\right)$, where $B\oplus Q\in M_{J}(2n-2,{\mathbb{R}})$ with $J=J_{1}\oplus J^{\prime}$. ∎
Let $A\in M(2n-1,{\mathbb{R}})$ be a non-nilpotent matrix. By combining Theorem 4.16, Theorem 4.17 and Lemma 3.9 it follows that $\mathfrak{g}_{A}$ admits a complex structure if and only if $A$ is conjugate to $bM$ for some $b\in{\mathbb{R}},\,b\neq 0$, where either $1\in\text{spec}\,(M)$ or $0\in\text{spec}\,(M)$ and $M$ takes one of the forms below:
  1. (1)
    $M=M(1)$ with $\Lambda(M(1))\in\bigwedge^{c}_{2n-1}$,
  2. (2)
    $M=M(1)\oplus Q$ where $\Lambda(M(1))\in\bigwedge^{c}_{m_{1}}$ with $m_{1}$ odd and $Q\in M_{J_{r}}(2r,{\mathbb{R}})$ with $2r=2n-1-m_{1}$ and $1\notin\text{spec}\,(Q)$,
  3. (3)
    $M=M(0)\oplus Q$ where $\Lambda(M(0))\in\bigwedge^{c}_{m_{0}}$ with $m_{0}$ odd and $Q\in M_{J_{r}}(2r,{\mathbb{R}})$ with $2r=2n-1-m_{0}$, $0\notin\text{spec}\,(Q)=\{\mu_{1},\ldots,\mu_{s}\}$ and $01=|\mu_{1}|\geq\cdots\geq|\mu_{s}|>$ 0.
We point out that, since $A$ is conjugate to $bM$ with $b\neq 0$, it follows from Proposition 2.1 that $\mathfrak{g}_{A}$ is isomorphic to $\mathfrak{g}_{M}$. Therefore, the isomorphism classes of $\mathfrak{g}_{A}$ are parametrized by the isomorphism classes of $\mathfrak{g}_{M}$. In order to obtain this parametrization, we define the following quotient spaces:
$\mathcal{C}^{1}_{r}:=\{Q\in M_{J_{r}}(2r,{\mathbb{R}})\,:\,1\notin\text{spec}\,(Q)\}/GL(2r,{\mathbb{R}}),$
$\mathcal{C}^{0}_{r}:=\{Q\in M_{J_{r}}(2r,{\mathbb{R}}):0\notin\text{spec}\,(Q)=\{\mu_{1},\ldots,\mu_{s}\}\text{ and }1=|\mu_{1}|\geq\cdots\geq|\mu_{s}|>0\}/_{\sim}$
where $Q_{1}\sim Q_{2}$ if and only if $Q_{1}$ is $GL(2r,{\mathbb{R}})$-conjugate to ±$\pm Q_{2}$ (see Lemma 3.9).
Corollary 4.18.
Let $A\in M(2n-1,{\mathbb{R}})$ be a non-nilpotent matrix. The isomorphism classes of $\mathfrak{g}_{A}$ admitting a complex structure are the isomorphism classes of $\mathfrak{g}_{M}$ for $M$ as in ${\rm(1)}$, ${\rm(2)}$ or ${\rm(3)}$ above, which are parametrized, respectively, by:
[…]
5. Classification of almost abelian Lie algebras admitting symplectic structures
5.1. Characterization of almost abelian Lie algebras admitting a symplectic structure